OpenStudy (anonymous):

Let f be the function with derivative given by f'(x) = sin(x^2 + 1). How many relative extrema does f have on the interval 2 < x < 4 ? How would I solve this without a calculator?

OpenStudy (anonymous):

i guess you have to find the zeros of \(f'\)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

not that hard since you know that \(\sin(0)=0\)

OpenStudy (anonymous):

But x^2 cannot be negative.

OpenStudy (anonymous):

oh good point, i thought it said \(\sin(x^2-1)\)

OpenStudy (anonymous):

That would be alot easier!

OpenStudy (anonymous):

might try setting \[x^2+1=\pi\] hten

OpenStudy (anonymous):

see if get something in the interval

OpenStudy (anonymous):

No, it's not on the interval.

OpenStudy (anonymous):

yes i think it is

OpenStudy (anonymous):

looks like there are two of them in fact http://www.wolframalpha.com/input/?i=sin%28x^2%2B1%29

OpenStudy (anonymous):

\[x^2+1=\pi\] \[x^2=\pi-1\] \[x=\pm\sqrt{\pi-1}\]

OpenStudy (anonymous):

That is 1.46 though.

OpenStudy (anonymous):

U WILL get one maxima and other minima x=3.4,x=3.8

OpenStudy (anonymous):

Yes, I know, there is a variable in front of the Root(pi-1), but just root(pi-1) did not fall in the interval. I have the answer, I needed an explination.

OpenStudy (anonymous):

\[\sin(x^2+1)=0 for \sin (npi)\] n=0,1,2,3,4 ......... g